# Explanation User Interfaces: A Systematic Literature Review

### v1

@@ 6.2. Tailoring XUI Evaluation Methods to User Groups
6.3. User group and Metrics
Figure 5. Reported evaluation criteria by user type. The stacked bar chart illustrates which user groups (AI experts, domain experts, non-experts, and unspecified) were associated with various evaluation criteria used in assessing explanation interfaces. Criteria such as trust, understandability, usability, and perceived effectiveness were evaluated across all user types, highlighting the multidimensional nature of user-centered evaluation in explainable systems. Notably, transparency was not evaluated in studies involving AI experts.
[…]
@@ 6.3. User group and Metrics
The relationship between the user type involved in the study and the metrics assessed provides insight into the key aspects that XUI designers prioritise when evaluating an application’s suitability for real-world adoption. Figure 5 reveals several key trends. Domain experts are the predominant group assessing helpfulness (Hwang et al., 2022a; Nourani et al., 2022b; Liu et al., 2022; Zytek et al., 2022); this might indicate that XUIs are designed with expert decision-making support in mind, where explainability is expected to enhance task performance rather than merely provide technical insights. Since domain experts are the end-users of AI systems in high-stakes applications (e.g., healthcare, finance), their evaluation of helpfulness is crucial in determining whether explanations effectively aid decision-making rather than just offering transparency.
Interestingly, AI experts are almost absent in studies assessing transparency, despite their technical expertise. This suggests that transparency is being evaluated primarily from a user-facing perspective rather than from an algorithmic interpretability standpoint. This metric frequently appears in the Conversational AI domain (Khurana et al., 2021; Hernandez-Bocanegra and Ziegler, 2023) and recommendation systems (Hernandez-Bocanegra and Ziegler, 2021; Guesmi et al., 2022; Lee et al., 2023; Park et al., 2022), highlighting its importance in ensuring domain experts and lay-users understand system decisions and interactions in these applications.
AI experts are more involved in studies that assess task performance, usability, and confidence: task performance is typically assessed in applications within computer science-specific domains, such as networking (Fujiwara et al., 2020; Xuan et al., 2022) and machine learning (Piorkowski et al., 2023). Usability studies involving AI experts span a broad range of domains, including mobility (Antar et al., 2024), education (Yuan et al., 2024b), economics (Purificato et al., 2023; Hohman et al., 2019), and agnostic applications (Cheng et al., 2021; Hao et al., 2023), reflecting a general need to ensure the usefulness of XUI applications. The confidence metric (Karran et al., 2022; Baniecki et al., 2024; Zhou et al., 2021), measured significantly less frequently compared to the others, indicates a concern for how well the system supports trust in its outputs, ensuring that explanations are not only technically accurate but also foster reliability and acceptance in real-world applications.
7. RQ4. What can guide the design of XUIs?
@@
Table 12 reports the list of papers that present design guidelines. We grouped the papers that propose a XUI according to the AI model used by the system. This categorization is meant to provide readers with a quick reference for the design of XUIs for systems that use a specific technology (i.e., Neural networks, Ensemble methods, Reinforcement Learning, Transparent Models, Generic AI models, Other models). Also, the list of papers that only present a framework for designing XUIs is reported.
@@
AI model References
Neural networks (Weitz et al., 2021; Hwang et al., 2022a; Fujiwara et al., 2020; Krause et al., 2017; Dominguez et al., 2020; Kim et al., 2023d; Wang et al., 2021; Lee et al., 2020; Huang et al., 2022; Malandri et al., 2023; Wang et al., 2023; Kim et al., 2023c; Nourani et al., 2022b; Meza Martínez et al., 2023; Hernandez-Bocanegra and Ziegler, 2021; Kim et al., 2023a; Guesmi et al., 2024; Lee et al., 2023; Wang et al., 2022; Esfahani et al., 2024; Lindvall et al., 2021; Zhang et al., 2024; Hao et al., 2023; Heimerl et al., 2022; Chotisarn et al., 2023; Gehrmann et al., 2019; Mishra et al., 2022; Ming et al., 2019; Karran et al., 2022)
Ensemble methods (Wysocki et al., 2023; Chromik et al., 2021; Wang et al., 2019; Cheng et al., 2022; Bhattacharya et al., 2024; Bove et al., 2023; Ma et al., 2022; Kalamaras et al., 2019; Cau et al., 2023)
Reinforcement Learning (Khanna et al., 2022; Park et al., 2022; Mishra et al., 2022)
Transparent models (Guo et al., 2022; Nakao et al., 2022; Bhattacharya et al., 2023; Hohman et al., 2019; Lee et al., 2020)
Generic/Model-agnostic models (Ferrario et al., 2020; Sanneman and Shah, 2022a; Cabour et al., 2023; Su et al., 2020; Naiseh et al., 2024; Mosca and Such, 2021; Fiok et al., 2022; Raees et al., 2024; Liu et al., 2022; Laato et al., 2022; Meyer and Zhu, 2024; Sanneman and Shah, 2022b; Ferreira and Monteiro, 2020; Simkute et al., 2022; Wu et al., 2022; Füßl et al., 2024; Khurana et al., 2021; Jansen et al., 2024; Cheng et al., 2021; Szymanski et al., 2024; Schneider et al., 2021a; Collaris and Van Wijk, 2020; Kim et al., 2023e; Naiseh et al., 2023; Okolo et al., 2024; Bahel et al., 2024; Larasati et al., 2023; Zytek et al., 2022; Szymanski et al., 2021; Lei et al., 2024)
Other AI models (Riveiro and Thill, 2021; Bove et al., 2022; Piorkowski et al., 2023; Desolda et al., 2023)
Frameworks only (Cabour et al., 2023; Su et al., 2020; Naiseh et al., 2024; Fiok et al., 2022; Meyer and Zhu, 2024; Sanneman and Shah, 2022b; Simkute et al., 2022; Raees et al., 2024; Liao et al., 2020; Pieters, 2011)
7.1. Core design principles for XUIs
@@
At the heart of any effective XUI is a design that is grounded in the user’s needs, context, and expertise. At a high level, explanations in XUIs should use clear, jargon-free language that adapts to the user’s context, reducing cognitive load and making AI behaviour more accessible (Khurana et al., 2021; Jansen et al., 2024; Okolo et al., 2024; Szymanski et al., 2024). User-centered frameworks such as (Barda et al., 2020) should be leveraged to define the purpose, content, and presentation of explanations in contexts such as clinical decision support (Hwang et al., 2022a). Ensemble-based XUIs emphasise the need to account for users’ mental models and to minimise cognitive load (Wysocki et al., 2023; Chromik et al., 2021), while generic systems advocate for tailoring explanations to both novice and expert users (Khurana et al., 2021; Jansen et al., 2024). Particular effort should be put in this aspect by researchers (e.g., using participatory design, conducting usability studies), as often solutions that appear effective to developers are not ideal to end-users (Guo et al., 2022).
@@
User trust is fostered by a high level of transparency of the AI’s decision-making process. To build trust, the potential and limitations of the AI system should be properly conveyed by ensuring that users receive accurate cues about model confidence and reliability (Okolo et al., 2024; Zytek et al., 2022). With black-box models such as neural networks, transparency can be achieved by displaying detailed insights, such as model uncertainties, confidence scores, and error metrics (Lee et al., 2023; Malandri et al., 2023; Nourani et al., 2022b; Zhang et al., 2024). Offering descriptive statistics and summaries of model performance can allow users to gauge overall system reliability (Lee et al., 2023; Malandri et al., 2023). White-box systems, in particular, stress that explanations should faithfully represent how the underlying model functions, including both local and global insights (Nakao et al., 2022; Guo et al., 2022), possibly to increase understanding of the risk behind a model’s decision (Bhattacharya et al., 2023). To increase trust, explanations should use an adequate language that is either accessible and avoids technical jargon (Okolo et al., 2024) or is tailored to the user’s domain and cultural context (Szymanski et al., 2024). Explanations should also resemble human reasoning and the way of explaining (Kim et al., 2023e). Finally, special care should also be taken in the selection of features to be addressed in explanations, as not every feature may be understandable and meaningful to the end user (Desolda et al., 2023).
@@
Placing explanations together with additional contextual information enhances their relevance and interpretability to domain-expert users, e.g., physicians (Hwang et al., 2022a; Cheng et al., 2022), AI engineers (Fujiwara et al., 2020; Bhattacharya et al., 2024), etc. Moreover, the depth of information in explanations should be adaptable to the user’s expertise, ensuring that both detailed and summary information are available without overwhelming the user (Wang et al., 2019; Naiseh et al., 2023; Lei et al., 2024).
@@
7.2. Frameworks
@@
Meyer and Zhu (Meyer and Zhu, 2024) examined 28 XUIs and linked specific UI components to explanation types. While “Why” and “What-if” questions are commonly addressed, tools for “How” or “Input” explanations are scarce, with elements like sliders, buttons, and icons playing key explanatory roles. Liao et al. (Liao et al., 2020) identified gaps between XAI research and design practice by interviewing 20 practitioners, resulting in an XAI question bank based on prototypical user inquiries. These questions are shaped by user context and mapped to suitable explanation techniques, such as global or counterfactual methods. Su et al. (Su et al., 2020) analyzed 40 mobile health apps and found that users want AI features—like recommendations or predictions—explained in clear, engaging ways. The study emphasises aligning explanations with user literacy and providing meaningful feedback loops to refine AI behavior.
@@
The Value Sensitive Design methodology (Naiseh et al., 2024) emphasises embedding human values—such as transparency, autonomy, and inclusivity—into explainable AI systems. Similarly, Raees et al. (Raees et al., 2024) advocate for user engagement and transparency to enhance trust in interactive AI. Sanneman and Shah (Sanneman and Shah, 2022b) proposed SAFE-AI, a framework for enhancing user situation awareness by addressing what AI did, why it did so, and what it will do. It also includes trust calibration and workload management to optimise user-AI interaction.
@@
In education, Fiok et al.(Fiok et al., 2022) advocate for participatory, user-centered design in XAI tools. Their work highlights the importance of grounding design in psychological theory and leveraging HCI expertise to align technical systems with learners’ needs and trust-building goals. Simkute et al.(Simkute et al., 2022) presented design principles for educational XUIs based on cognitive psychology; effective learning support includes techniques such as cognitive forcing, analogies, contrastive reasoning, and interactive feedback to promote skill development.
The Abstracted Explanation Space (Cabour et al., 2023) offers a structured framework that links user goals with appropriate explanation types, covering perception, comprehension, and projection. It advocates cross-disciplinary collaboration and helps designers tailor explanations to avoid cognitive overload and enhance relevance. Pieters (Pieters, 2011) distinguishes between explanation-for-confidence and explanation-for-trust, showing how explanation goals shape user understanding; the author argues that explanation depth must match its intent: too little detail may hinder trust, while too much may lower confidence.
8. The HERMES Platform to Guide Practitioners in Creating XUIs
@@
Figure 6. Distribution of explanation modalities across different application domains, data types, and types of users. The chart highlights how visual explanations dominate across most contexts, while other modalities—such as natural language, non-verbal, or tangible explanations—vary depending on the domain and target user group. This visualization reveals distinct preferences and trends in how explanations are tailored based on the nature of the data and the expertise of the intended users.

### v2

@@ 4.3.2. Tailoring XUI Evaluation Methods to User Groups
4.3.3. User group and Evaluation Constructs
Figure 6. Reported evaluation criteria by user type. Figure shows which user groups were associated with various evaluation criteria used in assessing explanation interfaces. Criteria such as trust, understandability, usability, and perceived effectiveness were evaluated across all user types, highlighting the multidimensional nature of user-centered evaluation in explainable systems. Notably, transparency was not evaluated in studies involving AI experts.
@@ 4.3.3. User group and Evaluation Constructs
The relationship between the user type involved in the study and the constructs assessed provides insight into the key aspects that XUI designers prioritise when evaluating an application’s suitability for real-world adoption. Figure 6 reveals several key trends. Domain experts are the predominant group assessing helpfulness (Hwang et al., 2022b; Nourani et al., 2022; Liu et al., 2022; Zytek et al., 2022); this might indicate that XUIs are designed with expert decision-making support in mind, where explainability is expected to enhance task performance rather than merely provide technical insights. Since domain experts are the end-users of AI systems in high-stakes applications (e.g., healthcare, finance), their evaluation of helpfulness is crucial in determining whether explanations effectively aid decision-making rather than just offering transparency. Interestingly, AI experts are almost absent in studies assessing transparency, despite their technical expertise. This suggests that transparency is being evaluated primarily from a user-facing perspective rather than from an algorithmic interpretability standpoint. This construct frequently appears in the Conversational AI domain (Khurana et al., 2021; Hernandez-Bocanegra and Ziegler, 2023) and recommendation systems (Hernandez-Bocanegra and Ziegler, 2021; Guesmi et al., 2022; Lee et al., 2023; Park et al., 2022), highlighting its importance in ensuring domain experts and lay-users understand system decisions and interactions in these applications. AI experts are more involved in studies that assess task performance, usability, and confidence: task performance is typically assessed in applications within computer science-specific domains, such as networking (Fujiwara et al., 2020; Xuan et al., 2022) and machine learning (Piorkowski et al., 2023). Usability studies involving AI experts span a broad range of domains, including mobility (Antar et al., 2024), education (Yuan et al., 2024b), economics (Purificato et al., 2023; Hohman et al., 2019), and agnostic applications (Cheng et al., 2021; Hao et al., 2023), reflecting a general need to ensure the usefulness of XUI applications. The confidence constructs (Karran et al., 2022; Baniecki et al., 2024; Zhou et al., 2021), measured significantly less frequently compared to the others, indicates a concern for how well the system supports trust in its outputs, ensuring that explanations are not only technically accurate but also foster reliability and acceptance in real-world applications.
4.4. RQ4. What can guide the design of XUIs?
@@
Table 12 reports the list of papers that present design guidelines. We grouped the papers that propose an XUI according to the AI model used by the system. This categorization is meant to provide readers with a quick reference for the design of XUIs for systems that use a specific technology (i.e., Neural networks, Ensemble methods, Reinforcement Learning, Transparent Models, Generic AI models, Other models). Also, the list of papers that only present a framework for designing XUIs is reported.
Summary of findings
XUIs should be designed using a user-centered approach, accommodating users’ mental models and favouring transparency and trust. Providing interactivity and control of explanations, providing multiple levels of visualizations, together with contextual information to better understand them, can help in this regard, as well as allowing users to personalize explanations based on their profile. Finally, popular frameworks can guide the design of XUIs by defining how explanations should be shaped (e.g., which questions they should answer), what values a XUI system should possess (e.g., transparency, engagement, control on interactivity), how users may be involved in the design process, and how to align user goals with system explanations.
@@
AI model Num. papers References
Neural networks 29 (Weitz et al., 2021; Hwang et al., 2022b; Fujiwara et al., 2020; Krause et al., 2017; Dominguez et al., 2020; Kim et al., 2023a; Wang et al., 2021; Lee et al., 2020; Huang et al., 2022; Malandri et al., 2023; Wang et al., 2023; Kim et al., 2023c; Nourani et al., 2022; Meza Martínez et al., 2023; Hernandez-Bocanegra and Ziegler, 2021; Kim et al., 2023d; Guesmi et al., 2024; Lee et al., 2023; Wang et al., 2022; Esfahani et al., 2024; Lindvall et al., 2021; Zhang et al., 2024; Hao et al., 2023; Heimerl et al., 2022; Chotisarn et al., 2023; Gehrmann et al., 2019; Mishra et al., 2022; Ming et al., 2019; Karran et al., 2022)
Ensemble methods 9 (Wysocki et al., 2023; Chromik et al., 2021; Wang et al., 2019; Cheng et al., 2022; Bhattacharya et al., 2024; Bove et al., 2023; Ma et al., 2022; Kalamaras et al., 2019; Cau et al., 2023)
Reinforcement Learning 3 (Khanna et al., 2022; Park et al., 2022; Mishra et al., 2022)
Transparent models 5 (Guo et al., 2022; Nakao et al., 2022; Bhattacharya et al., 2023; Hohman et al., 2019; Lee et al., 2020)
Generic/Model-agnostic models 30 (Ferrario et al., 2020; Sanneman and Shah, 2022a; Cabour et al., 2023; Su et al., 2020; Naiseh et al., 2024; Mosca and Such, 2021; Fiok et al., 2022; Raees et al., 2024; Liu et al., 2022; Laato et al., 2022; Meyer and Zhu, 2024; Sanneman and Shah, 2022b; Ferreira and Monteiro, 2020; Simkute et al., 2022; Wu et al., 2022; Füßl et al., 2024; Khurana et al., 2021; Jansen et al., 2024; Cheng et al., 2021; Szymanski et al., 2024; Schneider et al., 2021b; Collaris and Van Wijk, 2020; Kim et al., 2023e; Naiseh et al., 2023; Okolo et al., 2024; Bahel et al., 2024; Larasati et al., 2023; Zytek et al., 2022; Szymanski et al., 2021; Lei et al., 2024)
Other AI models 4 (Riveiro and Thill, 2021; Bove et al., 2022; Piorkowski et al., 2023; Desolda et al., 2023)
Frameworks only 10 (Cabour et al., 2023; Su et al., 2020; Naiseh et al., 2024; Fiok et al., 2022; Meyer and Zhu, 2024; Sanneman and Shah, 2022b; Simkute et al., 2022; Raees et al., 2024; Liao et al., 2020; Pieters, 2011)
4.4.1. Core design principles for XUIs
@@
At the heart of any effective XUI is a design that is grounded in the user’s needs, context, and expertise. At a high level, explanations in XUIs should use clear, jargon-free language that adapts to the user’s context, reducing cognitive load and making AI behaviour more accessible (Khurana et al., 2021; Jansen et al., 2024; Okolo et al., 2024; Szymanski et al., 2024). User-centered frameworks such as (Barda et al., 2020) should be leveraged to define the purpose, content, and presentation of explanations in contexts such as clinical decision support (Hwang et al., 2022b). Ensemble-based XUIs emphasise the need to account for users’ mental models and to minimise cognitive load (Wysocki et al., 2023; Chromik et al., 2021), while generic systems advocate for tailoring explanations to both novice and expert users (Khurana et al., 2021; Jansen et al., 2024). Particular effort should be put into this aspect by researchers (e.g., using participatory design, conducting usability studies), as often solutions that appear effective to developers are not ideal to end-users (Guo et al., 2022).
@@
User trust is fostered by a high level of transparency of the AI’s decision-making process. To build trust, the potential and limitations of the AI system should be properly conveyed by ensuring that users receive accurate cues about model confidence and reliability (Okolo et al., 2024; Zytek et al., 2022). With black-box models such as neural networks, transparency can be achieved by displaying detailed insights, such as model uncertainties, confidence scores, and error metrics (Lee et al., 2023; Malandri et al., 2023; Nourani et al., 2022; Zhang et al., 2024). Offering descriptive statistics and summaries of model performance can allow users to gauge overall system reliability (Lee et al., 2023; Malandri et al., 2023). White-box systems, in particular, stress that explanations should faithfully represent how the underlying model functions, including both local and global insights (Nakao et al., 2022; Guo et al., 2022), possibly to increase understanding of the risk behind a model’s decision (Bhattacharya et al., 2023). To increase trust, explanations should use an adequate language that is either accessible and avoids technical jargon (Okolo et al., 2024) or is tailored to the user’s domain and cultural context (Szymanski et al., 2024). Explanations should also resemble human reasoning and the way of explaining (Kim et al., 2023e). Finally, special care should also be taken in the selection of features to be addressed in explanations, as not every feature may be understandable and meaningful to the end user (Desolda et al., 2023).
@@
Placing explanations together with additional contextual information enhances their relevance and interpretability to domain-expert users, e.g., physicians (Hwang et al., 2022b; Cheng et al., 2022), AI engineers (Fujiwara et al., 2020; Bhattacharya et al., 2024), etc. Moreover, the depth of information in explanations should be adaptable to the user’s expertise, ensuring that both detailed and summary information are available without overwhelming the user (Wang et al., 2019; Naiseh et al., 2023; Lei et al., 2024).
@@
4.4.2. Frameworks
The ten frameworks resulting from the literature review cover heterogeneous aspects of the design of XUIs and may be grouped into three categories: (1) Interaction and Content Design Frameworks: These frameworks focus on the artifact (i.e., the explanation itself) and provide operational guidance on mapping user inquiries to specific interface components and explanation techniques. (2) Axiological and Relational Design Frameworks: These frameworks focus on human values and the relationship between the user and the AI, guiding the system’s high-level ethical positioning and prioritizing trust/confidence calibration, user engagement, and situational awareness. (3) Cognitive and Learning Alignment Frameworks: These frameworks focus on cognitive aspects of the user, drawing upon cognitive psychology and learning theories to align the XUI with human information processing capabilities, literacy levels, and learning goals.
Interaction and Content Design Frameworks
@@ 4.4.2. Frameworks
In (Meyer and Zhu, 2024), Meyer and Zhu examined 28 XUIs and linked specific UI components to explanation types. While “Why” and “What-if” questions are commonly addressed, tools for “How” or “Input” explanations are scarce, with elements like sliders, buttons, and icons playing key explanatory roles. Liao et al. (2020) identified gaps between XAI research and design practice by interviewing 20 practitioners, resulting in an XAI question bank based on prototypical user inquiries. These questions are shaped by user context and mapped to suitable explanation techniques, such as global or counterfactual methods. Su et al. (2020) analyzed 40 mobile health apps and found that users want AI features—like recommendations or predictions—explained in clear, engaging ways. The study emphasises aligning explanations with user literacy and providing meaningful feedback loops to refine AI behavior.
Axiological and Relational Design Frameworks
@@
The Value Sensitive Design methodology (Naiseh et al., 2024) emphasises embedding human values—such as transparency, autonomy, and inclusivity—into explainable AI systems. Similarly, Raees et al. (2024) advocate for user engagement and transparency to enhance trust in interactive AI. Sanneman and Shah (2022b) proposed SAFE-AI, a framework for enhancing user situation awareness by addressing what AI did, why it did so, and what it will do. It also includes trust calibration and workload management to optimise user-AI interaction. Pieters (2011) distinguishes between explanation-for-confidence and explanation-for-trust, showing how explanation goals shape user understanding; the author argues that explanation depth must match its intent: too little detail may hinder trust, while too much may lower confidence.
Cognitive and Pedagogical Alignment Frameworks
@@
In education, Fiok et al. (2022) advocate for participatory, user-centered design in XAI tools. Their work highlights the importance of grounding design in psychological theory and leveraging HCI expertise to align technical systems with learners’ needs and trust-building goals. Simkute et al. (2022) presented design principles for educational XUIs based on cognitive psychology; effective learning support includes techniques such as cognitive forcing, analogies, contrastive reasoning, and interactive feedback to promote skill development. The Abstracted Explanation Space (Cabour et al., 2023) offers a structured framework that links user goals with appropriate explanation types, covering perception, comprehension, and projection. This framework advocates cross-disciplinary collaboration and helps designers tailor explanations to avoid cognitive overload and enhance relevance.
5. Discussion
@@
Figure 7. Distribution of explanation modalities across different application domains, data types, and types of users. Visual explanations dominate across most contexts, while other modalities, such as natural language, non-verbal, or tangible explanations, vary depending on the domain and target user group. The visualization reveals distinct preferences and trends in how explanations are tailored based on the nature of the data and the expertise of the intended users.
This section discusses the results by relating the dimensions analyzed in our RQs. dditionally, to make our discussion actionable, we present a platform to support practitioners in building effective XUIs: HERMES. We examine how each dimension interacts across all RQs to identify recurring patterns. This approach allows us to distill insights that can support practitioners in navigating the broader landscape of the findings.
Visual and textual explanations are the primary means of conveying explanations, as shown in Figure 7, where the explanation modality is shown relative to its use within different domains, types of data, and types of users. This preference is particularly evident when working with tabular data and images, where graphical representations can make patterns and relationships more apparent. However, visual explanations are not limited to these data types; they are also applied to text and time series, while textual explanations remain common for tabular data. It is also interesting to note how in the mobility domain (Schneider et al., 2021a), audio is present.
The choice of explanation technique often depends on the audience. SHAP (Malandri et al., 2023; Ma et al., 2022; Chromik, 2021; He et al., 2024; Jansen et al., 2024) explanations tend to be preferred when working with non-expert users due to their intuitive visualization of feature contributions. However, exemplars i.e., concrete instances that illustrate a model’s behavior, and salient masks, which highlight the most influential input regions, are also commonly used to aid understanding (Huang et al., 2022; Wang and An, 2021; Yang et al., 2020; Larasati et al., 2023; Nakao et al., 2022). When designing user studies for XUIs, the choice of method should align with the target user group. Interviews are most commonly used for domain experts, particularly in high-stakes fields like healthcare and finance, where qualitative insights are crucial for understanding decision-making needs. In contrast, surveys and usability studies are more prevalent among AI experts, emphasizing practicality and efficiency in technical contexts. Regarding evaluation constructs, domain experts in high-stakes applications tend to prioritise helpfulness (Hwang et al., 2022a; Nourani et al., 2022; Liu et al., 2022; Zytek et al., 2022) over transparency (Sevastjanova et al., 2021). Their primary concern is whether an explanation supports their decision-making process rather than simply revealing model internals. For XUIs designers, this means that explanations should be crafted to disclose system behaviour and integrate seamlessly into expert workflows, ensuring they provide meaningful, actionable insights. The SLR presented in this paper has drawn on works from multiple disciplines to gather knowledge from diverse expertise fields. These vary substantially according to several factors, such as the application domain, the technologies underlying an AI system, and the type of users. Therefore, accessing these findings and integrating them into a specific design context would be cumbersome and require extensive knowledge of different fields (e.g., a designer should be knowledgeable about AI technologies). To make the findings more accessible, we developed a web platform called HERMES, which is described in the following.
5.1. Navigating the Results: Introducing the HERMES Platform